OpenStudy (ajanijones):

HELP please medals to. Use the distributive property to rewrite the expression and solve. 9*72 = 9* (blank) + 9*2 = (blank) Find the blank I think its 70 for the first one and 648 for the last?

OpenStudy (ajanijones):

1 second be right back.

OpenStudy (igreen):

Subtract..72 - 2 = ?

OpenStudy (ajanijones):

Thats 70.

OpenStudy (igreen):

\(9 \times 72 = 9 \times \_\_ + 9 \times 2 = \_\_\) Yes, so that's what goes in the first blank.

OpenStudy (ajanijones):

I already got 70 for the first blank spot and 648 for the last one.

OpenStudy (igreen):

Yep, you got it.

OpenStudy (texaschic101):

9 * 72 = 9 * 70 + 9 * 2 =

OpenStudy (ajanijones):

Thank,you guys.